OpenStudy (anonymous):

medal for whoever answers first!!!!!!!!!: a company distributes free pencils to all of the students in x schools. each school has (x+3) classes. the number of students in each class is 4 more than the number of classes in each school. each student is given (x+1) pencils. Part A: wright an expression to show the total number of pencils distributed by the company in x schools. Part B: what does x(x+3) represent? when simplified what would be the degree and classification? Part C:How can you calculate the total number of students in each school?

zepdrix (zepdrix):

Oh boy this question is so goofy.. Hmm let's see..

zepdrix (zepdrix):

Let's try to organize this: Number of schools: \(\Large\bf\sf x\) Number of classrooms: \(\Large\bf\sf (x+3)\) Number of students per classroom: \(\Large\bf\sf (x+3)+4\) Number of pencils per student: \(\Large\bf\sf (x+1)\)

OpenStudy (anonymous):

Exactly what I was doing lol, beat me to it

zepdrix (zepdrix):

Oh :3

OpenStudy (anonymous):

Pencils per classroom then would be: ((x+3)+4)(x+1)

OpenStudy (anonymous):

(x+3) classrooms, ((x+3)+4)(x+1) pencils per classroom. (x+3)[(x+3)+4](x+1) = pencils per school. Multiply by the number of schools, and you get: x(x+3)[(x+3)+4](x+1) = total pencils

OpenStudy (triciaal):

x(x+3)(x+1)(4 + x + 3)

OpenStudy (triciaal):

b) total number of classes

OpenStudy (triciaal):

x^2 + 3x degree is the highest power of the variable 2

OpenStudy (anonymous):

a) x(x+3)(x+7)(x+1) = x^4+11x^3+31x^2+21x. If you're unsure of which, I'd include both. b) (x+3) = classes per school, x = schools, so x(x+3) = total classes. x(x+3)=x^2+3x, degree 2 (greatest exponent). c) Chop off the pencils per student in the original equation (drop "(x+1)"). x(x+3)(x+7) or x^3+10x^2+21x

OpenStudy (anonymous):

okay so are you sure these are the correct answers?^^